### v1

# CdTe and HgTe doped with V, Cr, and Mn – prospects for the quantum anomalous Hall effect

Abstract
Using first principle calculations we examine properties of (Cd,V)Te, (Cd,Cr)Te, (Hg,V)Te, and (Hg,Cr)Te relevant to the quantum anomalous Hall effect (QAHE), such as the position of V- and Cr-derived energy levels and the exchange interactions between magnetic ions. We consider CdTe and HgTe, containing 12.5% of cation-substitutional V or Cr ions in comparison to the well-known case of (Cd,Mn)Te and (Hg,Mn)Te, and examine their suitability for fabrication of ferromagnetic barriers or ferromagnetic topological quantum wells, respectively. To account for the strong correlation of transition metal electrons we employ hybrid functionals with different mixing parameters focusing on , which better reproduces the experimental band gaps in HgTe and HgMnTe. We find that Cr, like Mn, acts as an isoelectronic dopant but V can be an in-gap donor in CdTe and a resonant donor in HgTe, similar to the case of Fe in HgSe. From the magnetic point of view, Cr-doping results in a ferromagnetic phase within the general gradient approximation (GGA) but interactions become antiferromagnetic within hybrid functionals. However, (Hg,V)Te is a ferromagnet within both exchange-correlation functionals in a stark contrast to (Hg,Mn)Te for which robust antiferromagnetic coupling is found theoretically and experimentally. Furthermore, we establish that the Jahn-Teller effect is relevant only in the case of Cr-doping. Considering lower defect concentrations in HgTe-based quantum wells compared to (Bi,Sb)Te layers, our results imply that HgTe quantum wells or (Cd,Hg)Te barriers containing either V or Cr show advantages over (Bi,Sb,Cr,V)Te-based QAHE systems but whether (i) ferromagnetic coupling will dominate in the Cr case and (ii) V will not introduce too many electrons to the quantum well is to be checked experimentally.
pacs:
71.15.-m, 71.15.Mb, 75.50.Cc, 74.40.Kb, 74.62.FjI Introduction
The theoretical prediction Yu et al. (2010) and the experimental discovery of the quantum anomalous Hall effect (QAHE) in dilute ferromagnetic semiconductor (Bi,Sb,Cr)Te Chang et al. (2013) and other systems Chang et al. (2023) have triggered research on the prospects of dissipationless and spin-polarized carrier channels for energy-efficient and decoherence-free electronic and spintronic classical and quantum devices Ke et al. (2018); Tokura et al. (2019). Simultaneously, the application potential of the QAHE for resistance Fox et al. (2018); Goetz et al. (2018); Yuma Okazaki et al. (2022) and current Rodenbach et al. (2023) standards operating in the absence of an external magnetic field has been demonstrated. It has, however, become clear that relatively large native defect concentrations in (Bi,Sb)Te and related systems, typically above cm, and the associated in-gap impurity-band charge transport, limit the standards’ operation to below 100 mK and 1 A Fox et al. (2018); Goetz et al. (2018); Yuma Okazaki et al. (2022); Rodenbach et al. (2023); Fijalkowski et al. (2021). Particularly relevant for the present work is the observation of the QAHE in (Bi,Sb)Te layers sandwiched between ferromagnetic (Zn,Cr)Te barriers Watanabe et al. (2019).
It is, therefore, interesting to consider the metrology prospects of HgTe and related systems, in which the native defect concentration is at the cm level Dietl (2023a, b). It was found that at the quantum well (QW) thickness corresponding to the topological phase transition, the quantum Hall (QH) plateau appears in weak magnetic fields and persists in a broad magnetic range of the magnetic fields König et al. (2007); Yahniuk et al. (2021), the observation relevant to the QHE metrology Yahniuk et al. (2021). The effect is particularly spectacular in the case of (Cd,Hg)Te/(Hg,Mn)Te QW, where the broad plateau begins at 50 mT Shamim et al. (2020). Those surprising observations were explained by an energetic overlap of the acceptor impurity band with the hole portion of the QW Dirac cone Dietl (2023a, b). The Coulomb gap, charge ordering, and the formation of bound magnetic polarons (in the Mn-doped samples) are the essential ingredients of the model Dietl (2023a, b). In the case of the studied samples Shamim et al. (2020), QHE dominates over a possible QAHE Liu et al. (2008), as the QAHE requires also the presence of a magnetic field – ferromagnetic coupling between Mn spins in II-VI compounds appears only if the hole density exceeds cm Dietl et al. (1997); Ferrand et al. (2000), whereas intrinsic Mn-Mn interactions are antiferromagnetic Mycielski et al. (1984).
In the present and companion paper Śliwa and Dietl (2023), we address the question of which cation-substitutional transition-metal (TM) impurities in barriers or QWs could lead to the QAHE in HgTe-based systems. The conditions to be considered experimentally and theoretically include:
1.
Sufficiently high – in a couple of percent range – solubility limits of particular dopants at the cation-substitutional positions and without aggregation. Importantly, the equilibrium limits can often be overcome by appropriate growth protocols Dietl et al. (2015) and, for instance, incorporation of Cr to ZnTe by low-temperature molecular beam epitaxy appears successful Watanabe et al. (2019).
2.
Ferromagnetic coupling between localized spins. We anticipate that Mn and Co dopants can be excluded, as antiferromagnetic superexchange dominates for these ions in II-VI compounds, also in the case of HgMnTe Mycielski et al. (1984). Similarly, Fe impurities do not appear appropriate, as they exhibit Van Vleck’s paramagnetism in HgTe Serre et al. (2006). Guided by the early theoretical prediction on the ferromagnetic superexchange in Cr-doped II-VI compounds Blinowski et al. (1996) as well by the observation of ferromagnetism at low temperatures in ZnCrTe Watanabe et al. (2019), we consider the Cr and V case.
3.
Sufficiently high Curie temperature at small enough in the case of QW doping to prevent a transition to the topologically trivial phase. In the case of HgMnTe, the inverted band structure disappears at % Sawicki et al. (1983) but QW confinement shifts to lower values Shamim et al. (2020). However, higher TM doping is possible in the case of barriers.
4.
Isoelectronic character of TM impurities, as charge doping by magnetic ions will hamper shifting of the Fermi level to the QW gap. The internal reference rule Langer et al. (1988), together with the known valence band offsets Mathieu et al. (1988); Dietl and Kossut (1988); Becker et al. (2000) and the positions of V levels in CdTe Selber et al. (1999) and Cr levels in CdTe Cieplak et al. (1975) and ZnTe Kuroda et al. (2007) suggest that the relevant TM donor level resides in the HgTe conduction and valence band, respectively. This indicates that V in either (Cd,Hg)Te barriers or HgTe QWs might act as an electron dopant, while Cr as an isoelectronic impurity.
5.
Formation of a single chiral edge channel in the presence of spin-polarized magnetic ions. According to the pioneering theoretical analysis Liu et al. (2008), the QAHE shows up in magnetically-doped HgTe quantum wells if , where and are the - and - exchange integrals, respectively. As discussed in the companion paper Śliwa and Dietl (2023), this condition can be relaxed by tilting the magnetization vector away from the growth direction, as in such a magnetization orientation, spin-orbit interaction diminishes spin-splitting of heavy-hole-like subbands and reduces an effective value.
6.
Single-ion magnetic anisotropy. Except for Mn for which orbital momentum is zero, magnetic ions such as V and Cr exhibit sizable single-ion anisotropy enlarged by the Jahn-Teller distortion. Accordingly, properties of HgTe QWs with those dopants are expected to be more sensitive to epitaxial strain than (Hg,Mn)Te QWs.
In this paper, we exploit a range of ab initio methods to asses three aspects of HgTe doped with 12.5% of cation-substitutional V and Cr ions. First, the concentration of V and Cr opens the band gap in (Hg,TM)Te. As mentioned, in the case of HgMnTe, the transition between the topological and non-topological phase occurs for but we find that can be larger for V and Cr.
Second, the positions of states brought about by V and Cr ions in CdTe and HgTe in comparison to the better-known case of Mn doping. Our results indicate that V impurities, in agreement with experimental results, act as mid-gap donors in CdTe but are close to the bottom of the conduction band in HgTe, so their isoelectronic character has to be checked experimentally. In contrast, Cr impurities do not provide carriers, as the relevant donor state is close to the valence band maximum in CdTe and deeper in the valence band of HgTe. However, compared to the Mn case, Cr donor level is much closer to the Fermi energy resulting in a relatively high magnitude of the - exchange integral compared to the (Hg,Mn)Te value.
Third, the magnitudes and signs of coupling between magnetic ions in those systems. We conclude that ferromagnetism dominates in the case of V ions, whereas there is a competition between ferromagnetic and antiferromagnetic coupling in the case of neighbor Cr spins.
II Computational details
In order to describe properly the band gap and strong correlation of electrons in the TM bands, we have used the Heyd-Scuseria-Ernzerhof 2006 (HSE06) hybrid functionalPaier et al. (2006). The mixing parameter is often set to around 0.20-0.30, and the value was used for Co-doped CdTe Dou et al. (2021). Other II-IV semiconductors have been studied with up to 0.36 Lyons et al. (2017); Kavanagh et al. (2021). We have carried out computations for , 0.32, and 0.5 with spin-orbit coupling (SOC) taken into account. Guided by our results presented in Sec. III and Appendix A, we focus on data for , which reproduces with accuracy of 0.1 eV experimental band gaps of CdTe, HgTe, CdMnTe, and HgMnTe. We have performed the band structure calculations using a plane-wave energy cutoff of 400 eV and an 888 -points grid centered in with 512 -points in the Brillouin zone, adopting the experimental lattice parameters, Å for HgTe and 6.4815 Å for CdTe Skauli and Colin (2001) and considering cation-substitutional transition metal content of 12.5 for all the compounds investigated. We put the spin polarization along the [001] direction. The band structure of undoped HgTe obtained with is shown in Fig. 1. For comparison, we have also performed computations within the GGA + approach. The obtained results are most similar to those obtained for .
Since V and Cr atoms are distributed periodically in supercells, the V- and Cr-derived states form bands. In reality, for randomly distributed magnetic ions, Anderson-Mott localization will result, at least at low magnetic ion concentrations, in strongly localized band gap levels or resonant states, which have a donor or acceptor character, and correspond to or states, respectively, where for V, Cr, and Mn, respectively. Furthermore, for the band structure determination, we arrange TM spins ferromagnetically along the [001] direction. Such a TM spins’ configuration leads to - exchange splitting of bands, which is -dependent due to the interplay of the - exchange with and spin-orbit interactionsAutieri et al. (2021).
III Band structures without distortions and prospects for QAHE
We start by showing electronic properties of the CdTe and HgTe doped with V and Cr without taking into account the distortions produced by the structural relaxation and assuming ferromagnetic ordering of TM spins. The band structures obtained within the HSE approach with for doped CdTe and HgTe with SOC interaction included and for ferromagnetic and periodic spin arrangement in comparison to the Mn-doped CdTe and HgTe are shown in Figs. 2 and 3, respectively. In contrast to the Mn case, where states are high in the conduction band and deep in the valence band, the states derived from V and Cr levels reside close to the bottom of the conduction band and the top of the valence band. In accord with experimental observations for CdTe:V Selber et al. (1999), the uppermost occupied -derived donor band in CdVTe lies in the middle of the band gap and the lowest unoccupied V-derived acceptor band near the bottom of the conduction band. In the case of CdCrTe, the occupied Cr-derived donor bands are strongly hybridized with the host valence band, whereas the Cr unoccupied acceptor band is near the bottom of the conduction band, again in accord with experimental data for CdTe:Cr Cieplak et al. (1975). A close proximity of the states to the top of the valence band results in a large magnitude of the - exchange integral and, therefore, in a larger splitting of the topmost valence band subbands near the point of the Brillouin zone in the Cr case compared to Mn doping, the effect clearly visible in Fig. 2. In accord with experimental results for HgMnTe, the band gap at the point in a paramagnetic phase, i.e., the distance between the centers of two split conduction subbands and four valence subbands implies normal band ordering with a gap of 0.2 eV. A similar band gap is implied by our results for HgCrTe meaning that QWs with Cr concentrations below 10% can show the QAHE. Furthermore, -dependent exchange splitting of bands leads to an overlap of the valence and conduction band states for certain directions in TM-doped HgTe, which points to the presence of outstanding carrier transport properties in and near the topological regime in those dilute magnetic systems.
Previous extensive studies of dilute magnetic materials Bonanni et al. (2021), together with the results for HgVTe presented in Fig. 3 for and also in Appendix A for and 0.5, indicate that four different scenarios relevant to the QAHE are possible:
1.
for certain V concentrations: (i) the band structure remains inverted and (ii) V donor states are in the valence and, therefore, V acts as an isoelectronic impurity, similarly to the case of Mn and, presumably Cr, in HgTe;
- 2.
3.
V acts as an electron dopant but does not give rise to the presence of resonant states;
4.
substantial hybridization between V orbitals and band states leads to unusual band ordering and minigaps in the vicinity of the Fermi energy.
In the case (1), (Hg,V)Te quantum wells can show the QAHE, as–according to our results presented in Sec. V–there is a ferromagnetic interaction between V spins in HgTe. The QAHE could be observed as long the QW is topological, i.e., neither a reduction in the QW thickness nor V doping makes the band ordering topologically trivial. By contrast, within scenarios (2) and (3), it will be difficult to shift the Fermi level from the conduction band to the topological gap for V concentrations sufficiently high to result in a ferromagnetic ground state. Finally, in the fourth case, presumably approximately described by ab initio results for HgVTe with ferromagnetic and periodic spin arrangement (Fig. 3), hybridization between host and dopant states leads to band reconstruction and hybridization gaps. It is unclear on whether the QAHE is possible under these conditions. Finally, we note that V in the (Cd,Hg,V)Te/HgTe quantum wells may act as a modulation electron dopant.
Another situation occurs in (Hg,Cr)Te QWs or (Cd,Hg,Cr)Te barriers, where the Cr donor-like states reside in the valence band but closer to the point compared to the Mn case. Accordingly, we expect Cr to be an isoelectronic dopant and to generate larger, compared to Mn doping, exchange splitting of subbands. These large splittings are seen in Fig. 3 which presents computational results for the ferromagnetic arrangement of TM spins along the [001] direction. These results substantiate also the tight-binding model of the (Hg,Cr)Te band structure employed to evaluate the sign and magnitude of exchange interactions between pairs of Cr impurities in HgTe Śliwa and Dietl (2023). As tight-binding and ab initio results presented in Sec. V point to a competition between antiferromagnetic and ferromagnetic interactions, the observation of the QAHE may require the application of a magnetic field in order to polarize Cr spins.
IV Local distortions produced by the Jahn-Teller effect
Lattice distortions produced by the Jahn-Teller (JT) effect in zinc-blende II-IV semiconductors are not straightforward. The impurities in ZnSeS were shown to produce octahedral rotations and changes of bond-angles Sarkar et al. (2022). Both Cr- and V-doping induce moderate tetrahedral distortions in CdTe and smaller tetrahedral distortions in HgTe. The main effect is a change in the Te-M-Te bond angle. In regular tetrahedra, we have four angles equal to 109.5 degrees. Among these four angles, two remain almost unchanged in the case of doping, one of these angles becomes smaller and another becomes larger. In Fig. 4 we show the crystal structure and the angles that change with the JT distortion.
| CdCrTe | CdVTe | HgCrTe | HgVTe | |
| Bond length (Å) | ||||
| M-Te | 2.8019 | 2.7622 | 2.8028 | 2.7282 |
| M-Te | 2.8040 | 2.7520 | 2.8057 | 2.7266 |
| M-Te | 2.7733 | 2.7600 | 2.7589 | 2.7271 |
| Angles (deg) | ||||
| 102.80 | 112.60 | 101.13 | 109.73 | |
| 115.72 | 107.64 | 117.07 | 109.25 | |
| PV(Å3) | 11.053 | 10.739 | 10.932 | 10.409 |
| BAV (deg) | 43.054 | 2.794 | 69.191 | 0.025 |
| DI (bond length) | 0.0040 | 0.0017 | 0.0062 | 0.0002 |
| (meV) | 34.696 | 4.400 | 45.05 | 2 20.094 |
When we dope the systems with Cr, the angle becomes larger and becomes smaller, the opposite happens when we dope with V. This situation is different from what happens in other systems Forte et al. (2018); Ni et al. (2020), where the two angles and become larger or smaller together. This allows a breaking of the degeneracy of the energy levels compatible with the cubic symmetry of the system. Two larger angles or two smaller angles will be more efficient in producing the JT splitting, and they will create a local tetragonal atom arrangement. In Table 1, we describe the local distortions in different systems and the gain in the total energy resulting from the JT distortion. These JT distortions help in stabilizing the localized character of states derived from TM levels.
V Exchange couplings between transition metal spins
In this section, we present computational results on exchange couplings between TM spins employing . In Table 2, we report energy differences between antiferromagnetic (AFM) and ferromagnetic (FM) configurations of TM spins in CdTe and HgTe doped with V, Cr, and Mn without taking the Jahn-Teller distortion into account. As seen, FM couplings prevail for the V case, whereas the interaction is AFM for Cr- and Mn-doped compounds. Once the Jahn-Teller distortion is taken into account, FM couplings show up in the Cr-compounds, too, as shown in Table 3. A comparison of data for presented here to results obtained for and 0.50 summarized in Appendix, demonstrate a strong sensitivity of the coupling strength to the employed theoretical framework. This sensitivity confirms the presence of a delicate balance between FM and AFM contributions to the TM coupling in the case of early transition metals in II-VI compounds Śliwa and Dietl (2023).
| compound | |||
| CdVTe | -22.3 | 78.3 | -3.4 |
| CdCrTe | -57.9 | -22.6 | -1.6 |
| CdMnTe | -16.2 | -16.1 | -0.3 |
| HgVTe | 35.5 | 41.2 | -2.7 |
| HgCrTe | -313.4 | -150.1 | -4.0 |
| HgMnTe | -17.0 | -16.9 | -1.6 |
| compound | ||
| CdCrTe NO JT | -57.9 | -22.6 |
| CdCrTe JT/2 | 10.4 | 4.8 |
| CdCrTe JT | 5.9 | 34.5 |
| HgCrTe NO JT | -313.4 | -150.1 |
| HgCrTe JT/2 | -100.4 | -154.4 |
| HgCrTe JT | -54.8 | -231.0 |
VI Conclusions
According to theoretical studies presented in the present and the companion paper Śliwa and Dietl (2023), the predicted magnitude and often sign of magnetic coupling between Cr ions in II-VI compounds depends on the approach employed. This fact reflects a competition between ferromagnetic superexchange and mostly antiferromagnetic Bloembergen-Rowland and two-electron terms Śliwa and Dietl (2023). Furthermore, according to results presented in Appendix in Fig. 6, pinning of the Fermi level below the top of the valence band by Cr ions in (Hg,Cr)Te cannot be entirely excluded. Ferromagnetic interactions appear more robust in V-doped compounds, though the donor character of V-impurities may show up in II-VI tellurides.
The ensemble of our results indicates that the most promising strategy for obtaining a QAHE device of HgTe-based systems is to employ (Cd,Hg,Cr)Te barriers that should be ferromagnetic already above 4.2 K, and in which - exchange splitting of bands should lead to splitting of QW states. That splitting can be further enhanced by a few percent Cr doping of the HgTe QW. However, the observation of the QAHE may require tilting of magnetization away from in-plane or perpendicular orientation Śliwa and Dietl (2023). Doping by V is expected to lead to a more robust ferromagnetic ground state but it could be that the resulting electron doping may preclude the shifting of the Fermi level to the topological gap. In any case, a close energetic proximity of states and the Fermi level opens doors for a new physics in both (Hg,V)Te and (Hg,Cr)Te, not encountered in the case of (Hg,Mn)Te.
Acknowledgments
The work is supported by the Foundation for Polish Science through the International Research Agendas program co-financed by the European Union within the Smart Growth Operational Programme (Grant No. MAB/2017/1). We acknowledge the access to the computing facilities of the Interdisciplinary Center of Modeling at the University of Warsaw, Grant G84-0, GB84-1, and GB84-7. We acknowledge the CINECA award under the ISCRA initiative IsC85 ”TOPMOST” and IsC93 ”RATIO” grant, for the availability of high-performance computing resources and support. We acknowledge the access to the computing facilities of the Poznan Supercomputing and Networking Center Grant No. 609.
Appendix A Transition metal doped compounds with 25% and 50% of the exact exchange
In this Appendix, we present band structures and the magnetic properties for TM-doped CdTe and HgTe obtained employing the hybrid functional with and of the exact exchange. The band structure results displayed in Figs. 5, 6, and 7 substantiate the use of in the main body of the text. In particular, in the case of HgMnTe, the experimental value of the band gap eV Furdyna (1988) is close to the computational data for (Fig. 3) but is too small and too large for and 0.5, respectively, as shown in Figs. 6 and 7.
As we can see in Fig. 6, for we find a metallic phase in the case of HgTe doped with Cr because acceptor -states of the dopants undergo a shift below the top of the valence band. The band structures obtained within the GGA+ approach resemble the results obtained at for all compounds and, therefore, we do not display them.
In Tables 4 and Table 5 we report the values of the spin-spin couplings for the Cr-doped compounds obtained by using and 0.5, respectively. In the latter case, we find that HgTe doped with Cr shows ferromagnetic couplings.
| compound | |||
| CdVTe | -38.1 | 22.8 | 0.6 |
| CdCrTe | -10.5* | -16.8* | -4.8 |
| CdMnTe | -21.9 | -21.8 | -0.5 |
| HgVTe | 47.6 | 51.5 | -1.6 |
| HgCrTe | -77.9 | -35.9 | -5.1 |
| HgMnTe | -22.3 | -22.2 | -2.0 |
| compound | ||
| CdCrTe NO JT | -218.7 | -160.3 |
| CdCrTe JT/2 | -62.3 | 24.1 |
| HgCrTe NO JT | 74.0 | 7.8 |
| HgCrTe JT/2 | 88.6 | -22.7 |