### v2

# CdTe and HgTe doped with V, Cr, and Mn – prospects for the quantum anomalous Hall effect

Abstract
Using first principle calculations we examine properties of (Cd,V)Te, (Cd,Cr)Te, (Hg,V)Te, and (Hg,Cr)Te relevant to the quantum anomalous Hall effect (QAHE), such as the position of V- and Cr-derived energy levels and the exchange interactions between magnetic ions. We consider CdTe and HgTe, containing 12.5% of cation-substitutional V or Cr ions in comparison to the well-known case of (Cd,Mn)Te and (Hg,Mn)Te, and examine their suitability for fabrication of ferromagnetic barriers or ferromagnetic topological quantum wells, respectively. To account for the strong correlation of transition metal electrons we employ hybrid functionals with different mixing parameters focusing on , which better reproduces the experimental band gaps in HgTe and Hg0.875Mn0.125Te. We find that Cr, like Mn, acts as an isoelectronic dopant but V can be an in-gap donor in CdTe and a resonant donor in HgTe, similar to the case of Fe in HgSe. From a magnetic point of view, the Cr-doped HgTe is ferromagnetic within the general gradient approximation (GGA) but becomes AFM within hybrid functionals. However, (Hg,V)Te is a ferromagnet within both exchange-correlation functionals in a stark contrast to (Hg,Mn)Te for which robust antiferromagnetic coupling is found theoretically and experimentally. Furthermore, we establish that the Jahn-Teller effect is relevant for the magnetism in the case of Cr-doping. Considering lower defect concentrations in HgTe-based quantum wells compared to (Bi,Sb)3Te2 layers, our results imply that HgTe quantum wells or (Cd,Hg)Te barriers containing either V or Cr show advantages over (Bi,Sb,Cr,V)3Te2-based QAHE systems but whether (i) ferromagnetic coupling will dominate in the Cr case and (ii) V will not introduce too many electrons to the quantum well is to be checked experimentally.
pacs:
71.15.-m, 71.15.Mb, 75.50.Cc, 74.40.Kb, 74.62.FjI Introduction
The theoretical prediction Yu et al. (2010) and the experimental discovery of the quantum anomalous Hall effect (QAHE) in dilute ferromagnetic semiconductor (Bi,Sb,Cr)2Te3 Chang et al. (2013) and other systems Chang et al. (2023) have triggered research on the prospects of dissipationless and spin-polarized carrier channels for energy-efficient and decoherence-free electronic and spintronic classical and quantum devices Ke et al. (2018); Tokura et al. (2019). Simultaneously, the application potential of the QAHE for resistance Fox et al. (2018); Goetz et al. (2018); Yuma Okazaki et al. (2022) and current Rodenbach et al. (2023) standards operating in the absence of an external magnetic field has been demonstrated. It has, however, become clear that relatively large native defect concentrations in (Bi,Sb)2Te3 and related systems, typically above cm-3, and the associated in-gap impurity-band charge transport, limit the standards’ operation to below 100 mK and 1 A Fox et al. (2018); Goetz et al. (2018); Yuma Okazaki et al. (2022); Rodenbach et al. (2023); Fijalkowski et al. (2021). Particularly relevant for the present work is the observation of the QAHE in (Bi,Sb)2Te3 layers sandwiched between ferromagnetic (Zn,Cr)Te barriers Watanabe et al. (2019).
It is, therefore, interesting to consider the metrology prospects of HgTe and related systems, in which the native defect concentration is at the cm-3 level Dietl (2023a, b). It was found that at the quantum well (QW) thickness corresponding to the topological phase transition, the quantum Hall (QH) plateau appears in weak magnetic fields and persists in a broad magnetic range of the magnetic fields König et al. (2007); Yahniuk et al. (2021), the observation relevant to the QHE metrology Yahniuk et al. (2021). The effect is particularly spectacular in the case of (Cd,Hg)Te/(Hg,Mn)Te QW, where the broad plateau begins at 50 mT Shamim et al. (2020). Those surprising observations were explained by an energetic overlap of the acceptor impurity band with the hole portion of the QW Dirac cone Dietl (2023a, b). The Coulomb gap, charge ordering, and the formation of bound magnetic polarons (in the Mn-doped samples) are the essential ingredients of the model Dietl (2023a, b). In the case of the studied samples Shamim et al. (2020), QHE dominates over a possible QAHE Liu et al. (2008), as the QAHE requires also the presence of a magnetic field – ferromagnetic coupling between Mn spins in II-VI compounds appears only if the hole density exceeds cm-3 Dietl et al. (1997); Ferrand et al. (2000), whereas intrinsic Mn-Mn interactions are antiferromagnetic Mycielski et al. (1984).
In the present and companion paper Śliwa and Dietl (2023), we address the question of which cation-substitutional transition-metal (TM) impurities in barriers or QWs could lead to the QAHE in HgTe-based systems. The conditions to be considered experimentally and theoretically include:
1.
Sufficiently high – in a couple of percent range – solubility limits of particular dopants at the cation-substitutional positions and without aggregation. Importantly, the equilibrium limits can often be overcome by appropriate growth protocols Dietl et al. (2015) and, for instance, incorporation of Cr to ZnTe by low-temperature molecular beam epitaxy appears successful Watanabe et al. (2019).
2.
Ferromagnetic coupling between localized spins. We anticipate that Mn and Co dopants can be excluded, as antiferromagnetic superexchange dominates for these ions in II-VI compounds, also in the case of Hg1-xMnxTe Mycielski et al. (1984). Similarly, Fe impurities do not appear appropriate, as they exhibit Van Vleck’s paramagnetism in HgTe Serre et al. (2006). Guided by the early theoretical prediction on the ferromagnetic superexchange in Cr-doped II-VI compounds Blinowski et al. (1996) as well by the observation of ferromagnetism at low temperatures in Zn1-xCrxTe Watanabe et al. (2019), we consider the Cr and V case.
3.
Sufficiently high Curie temperature at small enough in the case of QW doping to prevent a transition to the topologically trivial phase. In the case of Hg1-xMnxTe, the inverted band structure disappears at % Sawicki et al. (1983) but QW confinement shifts to lower values Shamim et al. (2020). However, higher TM doping is possible in the case of barriers.
4.
Isoelectronic character of TM impurities, as charge doping by magnetic ions will hamper shifting of the Fermi level to the QW gap. The internal reference rule Langer et al. (1988), together with the known valence band offsets Mathieu et al. (1988); Dietl and Kossut (1988); Becker et al. (2000) and the positions of V levels in CdTe Selber et al. (1999) and Cr levels in CdTe Cieplak et al. (1975) and ZnTe Kuroda et al. (2007) suggest that the relevant TM2+/3+ donor level resides in the HgTe conduction and valence band, respectively. This indicates that V in either (Cd,Hg)Te barriers or HgTe QWs might act as an electron dopant, while Cr as an isoelectronic impurity.
5.
Formation of a single chiral edge channel in the presence of spin-polarized magnetic ions. According to the pioneering theoretical analysis Liu et al. (2008), the QAHE shows up in magnetically-doped HgTe quantum wells if , where and are the - and - exchange integrals, respectively. As discussed in the companion paper Śliwa and Dietl (2023), this condition can be relaxed by tilting the magnetization vector away from the growth direction, as in such a magnetization orientation, spin-orbit interaction diminishes spin-splitting of heavy-hole-like subbands and reduces an effective value.
6.
Single-ion magnetic anisotropy. Except for Mn2+ for which orbital momentum is zero, magnetic ions such as V2+ and Cr2+ exhibit sizable single-ion anisotropy enlarged by the Jahn-Teller distortion. Accordingly, properties of HgTe QWs with those dopants are expected to be more sensitive to epitaxial strain than (Hg,Mn)Te QWs.
Our ab initio approach to the issues listed above is built on the previous extensive first-principles investigations of non-topological semiconductors Sato et al. (2010); Dietl et al. (2015) and topological Bi-Sb chalcogenides doped with transition metals Zhang et al. (2012); Vergniory et al. (2014); Bouaziz et al. (2018), as well as on results obtained recently for a wide family of Eu compounds, such EuCd2As2 and EuCd2Bi2 Cuono et al. (2023). Those studies have demonstrated that strong correlations on localized magnetic ions and narrow band gaps specific to topological systems require a careful selection and testing of the density-functional theory (DFT) implementation.
In this paper, we exploit a range of ab initio methods to asses three aspects of HgTe doped with 12.5% of cation-substitutional V and Cr ions. First, the concentration of V and Cr opens the band gap in (Hg,TM)Te. As mentioned, in the case of Hg1-xMnxTe, the transition between the topological and non-topological phase occurs for but we find that can be larger for V and Cr.
Second, the positions of states brought about by V and Cr ions in CdTe and HgTe in comparison to the better-known case of Mn doping. Our results indicate that V impurities, in agreement with experimental results, act as mid-gap donors in CdTe but are close to the bottom of the conduction band in HgTe, so their isoelectronic character has to be checked experimentally. In contrast, Cr impurities do not provide carriers, as the relevant donor state is close to the valence band maximum in CdTe and deeper in the valence band of HgTe. However, compared to the Mn case, Cr donor level is much closer to the Fermi energy resulting in a relatively high magnitude of the - exchange integral compared to the (Hg,Mn)Te value.
Third, the magnitudes and signs of coupling between magnetic ions in those systems. We conclude that ferromagnetism dominates in the case of V ions, whereas there is a competition between ferromagnetic and antiferromagnetic coupling in the case of neighbor Cr spins.
II Computational details
In order to describe properly the band gap and strong correlation of electrons in the TM bands, we have used the Heyd-Scuseria-Ernzerhof 2006 (HSE06) hybrid functionalPaier et al. (2006). The mixing parameter is often set to around 0.20-0.30, and the value was used for Co-doped CdTe Dou et al. (2021). Other II-IV semiconductors have been studied with up to 0.36 Lyons et al. (2017); Kavanagh et al. (2021). We have carried out computations for , 0.32, and 0.5 with spin-orbit coupling (SOC) taken into account. Guided by our results presented in Sec. III and Appendix A, we focus on data for , which reproduces with accuracy of 0.1 eV experimental band gaps of CdTe, HgTe, Cd0.875Mn0.125Te, and Hg0.875Mn0.125Te. We have performed the band structure calculations using a plane-wave energy cutoff of 400 eV and an 888 -points grid centered in with 512 -points in the Brillouin zone, adopting the experimental lattice parameters, Å for HgTe and 6.4815 Å for CdTe Skauli and Colin (2001) and considering cation-substitutional transition metal content of 12.5 for all the compounds investigated. We put the spin polarization along the [001] direction. The band structure of undoped HgTe obtained with is shown in Fig. 1. For comparison, we have also performed computations within the GGA + approach. The obtained results are most similar to those obtained for .
Since V and Cr atoms are distributed periodically in supercells, the V- and Cr-derived states form bands. In reality, for randomly distributed magnetic ions, Anderson-Mott localization will result, at least at low magnetic ion concentrations, in strongly localized band gap levels or resonant states, which have a donor or acceptor character, and correspond to or states, respectively, where for V, Cr, and Mn, respectively. Furthermore, for the band structure determination, we arrange TM spins ferromagnetically along the [001] direction. Such a TM spins’ configuration leads to - exchange splitting of bands, which is -dependent due to the interplay of the - exchange with and spin-orbit interactionsAutieri et al. (2021).
III Band structures without distortions and prospects for QAHE
We start by showing electronic properties of the CdTe and HgTe doped with V and Cr without taking into account the distortions produced by the structural relaxation and assuming ferromagnetic ordering of TM spins. The band structures obtained within the HSE approach with for doped CdTe and HgTe with SOC interaction included and for ferromagnetic and periodic spin arrangement in comparison to the Mn-doped CdTe and HgTe are shown in Figs. 2 and 3, respectively. In contrast to the Mn case, where states are high in the conduction band and deep in the valence band, the states derived from V and Cr levels reside close to the bottom of the conduction band and the top of the valence band. In accord with experimental observations for CdTe:V Selber et al. (1999), the uppermost occupied -derived donor band in Cd0.875V0.125Te lies in the middle of the band gap and the lowest unoccupied V-derived acceptor band near the bottom of the conduction band. In the case of Cd0.875Cr0.125Te, the occupied Cr-derived donor bands are strongly hybridized with the host valence band, whereas the Cr unoccupied acceptor band is near the bottom of the conduction band, again in accord with experimental data for CdTe:Cr Cieplak et al. (1975). In Fig. 4, a zoom of the bands close to the Fermi level is reported for HgTe and for HgTe doped with V, Cr, and Mn assuming a ferromagnetic ordering of the TM spins. The exchange splitting of the bands into six components corresponding to (sz = 1/2) and (jz = 3/2, 1/2) levels, is clearly seen. Interestingly, the subband arrangement and the magnitude of splitting differ for particular magnetic ions. In particular, the splitting of the band is seen to be larger in (Hg,V)Te and (Hg,Cr)Te compared to the (Hg,Mn)Te. We assign this difference to the proximity of the V and Cr levels to the top of the valence band, which enhances the - hybridization and the - Schrieffer-Wolff kinetic exchange, as discussed in detail in the companion paper Śliwa and Dietl (2023). At the same time, the - splitting of the band is smaller and roughly independent of the TM ion, as could be expected for the potential exchange. Another relevant information concerns the relative position of the and bands. In the case of HgTe, the band structure is inverted, i. e., the band is below the bands, which accounts for topological properties of HgTe and its quantum structures. According to results in Fig. 4, Cr and Mn doping shifts the subbands up with respect to the subbands, the fact already known experimentally for (Hg,Mn)Te Furdyna (1988) and (Hg,Mn)Te QWs Shamim et al. (2020), meaning that QWs with Cr content below 10 can show topological properties. Furthermore, -dependent exchange splitting of bands leads to an overlap of the valence and conduction band states for certain directions in TM-doped HgTe, which points to outstanding carrier transport properties in and near the topological regime in those dilute magnetic systems. Prospects, in the light of our results, of HgTe doped with TM ions for studies and applications of the quantum anomalous Hall effect are discussed in Sec. VI.
A close proximity of the states to the top of the valence band results in a large magnitude of the - exchange integral and, therefore, in a larger splitting of the topmost valence band subbands near the point of the Brillouin zone in the Cr case compared to Mn doping, the effect clearly visible in Fig. 2. In Fig. 4 a zoom of the bands of HgTe doped with V and Cr and close to the Fermi level is reported. The splitting of the bands is shown. There is a band inversion between the and the bands of the spin-up subsector. The topological invariant, which is the Chern number in this case, is extremely challenging to compute within the DFT approach for a strongly-correlated system with a supercell due to the doping. However, the detailed calculation of the Chern number was performed for the same system within the tight-binding model in the companion paperŚliwa and Dietl (2023) reporting the expected value of =1. The energy difference between and is 46.6 meV for Hg0.875V0.125Te and 47.9 meV for Hg0.875Cr0.125Te. In accord with experimental results for Hg1-xMnxTe, the band gap at the point in a paramagnetic phase, i.e., the distance between the centers of two split conduction subbands and four valence subbands implies normal band ordering with a gap of 0.2 eV. A similar band gap is implied by our results for Hg0.875Cr0.125Te meaning that QWs with Cr concentrations below 10% can show the QAHE. Furthermore, -dependent exchange splitting of bands leads to an overlap of the valence and conduction band states for certain directions in TM-doped HgTe, which points to the presence of outstanding carrier transport properties in and near the topological regime in those dilute magnetic systems.
| Cd1-xVxTe | Cd1-xCrxTe | Hg1-xVxTe | Hg1-xCrxTe | |
| Bond length (Å) | ||||
| M-Te1 | 2.8019 | 2.7622 | 2.8028 | 2.7282 |
| M-Te2 | 2.8040 | 2.7520 | 2.8057 | 2.7266 |
| M-Te3 | 2.7733 | 2.7600 | 2.7589 | 2.7271 |
| Angles (deg) | ||||
| 102.80 | 112.60 | 101.13 | 109.73 | |
| 115.72 | 107.64 | 117.07 | 109.25 | |
| PV(Å3) | 11.053 | 10.739 | 10.932 | 10.409 |
| BAV (deg2) | 43.054 | 2.794 | 69.191 | 0.025 |
| DI (bond length) | 0.0040 | 0.0017 | 0.0062 | 0.0002 |
| (meV) | 34.696 | 4.400 | 45.05 | 20.094 |
Previous extensive studies of dilute magnetic materials Bonanni et al. (2021), together with the results for Hg0.815V0.125Te presented in Fig. 3 for and also in Appendix A for and 0.5, indicate that four different scenarios relevant to the QAHE are possible:
1.
for certain V concentrations: (i) the band structure remains inverted and (ii) V donor states are in the valence and, therefore, V acts as an isoelectronic impurity, similarly to the case of Mn and, presumably Cr, in HgTe;
- 2.
3.
V acts as an electron dopant but does not give rise to the presence of resonant states;
4.
substantial hybridization between V orbitals and band states leads to unusual band ordering and minigaps in the vicinity of the Fermi energy.
In the case (1), (Hg,V)Te quantum wells can show the QAHE, as–according to our results presented in Sec. V–there is a ferromagnetic interaction between V spins in HgTe. The QAHE could be observed as long the QW is topological, i.e., neither a reduction in the QW thickness nor V doping makes the band ordering topologically trivial. By contrast, within scenarios (2) and (3), it will be difficult to shift the Fermi level from the conduction band to the topological gap for V concentrations sufficiently high to result in a ferromagnetic ground state. Finally, in the fourth case, presumably approximately described by ab initio results for Hg0.815V0.125Te with ferromagnetic and periodic spin arrangement (Fig. 3), hybridization between host and dopant states leads to band reconstruction and hybridization gaps. It is unclear on whether the QAHE is possible under these conditions. Finally, we note that V in the (Cd,Hg,V)Te/HgTe quantum wells may act as a modulation electron dopant.
Another situation occurs in (Hg,Cr)Te QWs or (Cd,Hg,Cr)Te barriers, where the Cr donor-like states reside in the valence band but closer to the point compared to the Mn case. Accordingly, we expect Cr to be an isoelectronic dopant and to generate larger, compared to Mn doping, exchange splitting of subbands. These large splittings are seen in Fig. 3 which presents computational results for the ferromagnetic arrangement of TM spins along the [001] direction. These results substantiate also the tight-binding model of the (Hg,Cr)Te band structure employed to evaluate the sign and magnitude of exchange interactions between pairs of Cr impurities in HgTe Śliwa and Dietl (2023). As tight-binding presented in the companion paperŚliwa and Dietl (2023) and ab initio results presented in Sec. V point to a competition between antiferromagnetic and ferromagnetic interactions, the observation of the QAHE may require the application of a magnetic field in order to polarize Cr spins.
IV Local distortions produced by the Jahn-Teller effect
While the long-range cooperative Jahn-Teller (JT) effect in homogeneous bulk systems can be described with a finite number of phonon modes as in molecules, the description of the JT effect in solids with JT-active defects involves an infinite number of phonon modes. A particularly tricky case is represented by the case of JT-active impurities in tetrahedra environments like CdTe and HgTe which are analyzed in this paper. Recently, Cruz et al.Cruz et al. (2024) presented a recipe to tackle this problem in the case of topologically trivial CdSe which is similar to CdTe, while we are not aware of attempting to solve the same problem in topologically non-trivial systems like HgTe. Therefore, the lattice distortions produced by the JT effect in zinc-blende II-IV semiconductors are not straightforward and the effect can be relevant. For instance, the impurities in zinc-blende ZnSexS1-x were shown to produce octahedral rotations and changes of bond-angles Sarkar et al. (2022). The JT lifted the degeneracy of the t2g and eg manifold for both spins in both Cr and V doped systems, this was established through the check of the electronic levels at the point. The lifting of the degeneracy of the manifold is necessary to observe an insulating phase. The main effect is a change in the Te-M-Te bond angle. In regular tetrahedra, we have four angles equal to 109.5 degrees. Among these four angles, two remain almost unchanged in the case of doping, one of these angles becomes smaller and another becomes larger. In Fig. 5 we show the crystal structure and the angles that change with the JT distortion. The bond angle variance (BAV) is much larger for the V-doping than for the Cr-doping. While the eg manifold for the majority spin is full, the V-doping consists of 1 electron in the t2g manifold, while the Cr-doping consists of 2 electrons in the t2g manifold. We assume that the BAV is larger for the case of a single electron in the manifold as happens for the d4 in an octahedral crystal field.
When we dope the systems with Cr, the angle becomes larger and becomes smaller, the opposite happens when we dope with V. This situation is different from what happens in other systems Forte et al. (2018); Ni et al. (2020), where the two angles and become larger or smaller together. The JT allows the breaking of the degeneracy of the energy levels compatible with the cubic symmetry of the system. Two larger angles or two smaller angles will be more efficient in producing the JT splitting, and they will create a local tetragonal atom arrangement. In Table 1, we describe the local distortions in different systems and the gain in the total energy resulting from the JT distortion. These JT distortions help in stabilizing the localized character of states derived from TM levels.
V Exchange couplings between transition metal spins
In this Section, we present computational results on exchange couplings between TM spins employing . In Table 2, we report energy differences between antiferromagnetic (AFM) and ferromagnetic (FM) configurations of TM spins in CdTe and HgTe doped with V, Cr, and Mn without taking the Jahn-Teller distortion into account. As seen, FM couplings prevail for the V case, whereas the interaction is AFM for Cr- and Mn-doped compounds. Once the Jahn-Teller distortion is taken into account, FM couplings show up in the Cr-compounds, too, as shown in Table 3. Regarding the V-doped systems, the ground state is still FM if we include the Jahn-Teller distortions, as shown in Table 4. A comparison of data for presented here to results obtained for and 0.50 summarized in Appendix, demonstrate a strong sensitivity of the coupling strength to the employed theoretical framework. This sensitivity confirms the presence of a delicate balance between FM and AFM contributions to the TM coupling in the case of early transition metals in II-VI compounds Śliwa and Dietl (2023).
| compound | |||
| Cd1-xVxTe | -22.3 | 78.3 | -3.4 |
| Cd1-xCrxTe | -57.9 | -22.6 | -1.6 |
| Cd1-xMnxTe | -16.2 | -16.1 | -0.3 |
| Hg1-xVxTe | 35.5 | 41.2 | -2.7 |
| Hg1-xCrxTe | -313.4 | -150.1 | -4.0 |
| Hg1-xMnxTe | -17.0 | -16.9 | -1.6 |
| compound | |||
| Cd1-xCrxTe NO JT | -57.9 | -22.6 | -1.6 |
| Cd1-xCrxTe JT/2 | 10.4 | 4.8 | -1.6 |
| Cd1-xCrxTe JT | 5.9 | 34.5 | -1.7 |
| Hg1-xCrxTe NO JT | -313.4 | -150.1 | -4.0 |
| Hg1-xCrxTe JT/2 | -100.4 | -154.4 | -4.6 |
| Hg1-xCrxTe JT | -54.8 | -231.0 | -5.6 |
| compound | |||
| Cd1-xVxTe NO JT | -22.3 | 78.3 | -3.4 |
| Cd1-xVxTe JT | 28.3 | 28.4 | -3.1 |
| Hg1-xVxTe NO JT | 35.5 | 41.2 | -2.7 |
| Hg1-xVxTe JT | 0.2 | 30.5 | -7.5 |
VI Implications to the anomalous Hall effect
Quantum wells of HgTe constitute a rare example of systems showing the quantum spin Hall effect. Our studies presented here and in the companion paper Śliwa and Dietl (2023) were designed to address the question about the suitability of HgTe QWs and related systems for the demonstration of the quantum anomalous Hall effect (QAHE) with characteristics opening a door or applications in the quantum metrology. As mentioned in Introduction, this question is timely considering the fact that Hall resistance quantization in (Bi,Sb,Cr,V) layers is perturbed above 50 mK by hopping conductivity Kawamura et al. (2017) involving numerous in-gap states brought about by antisites and other defects Zhang et al. (2012). As already noted by Liu et al. Liu et al. (2008), the appearance of the chiral edge states requires the presence of: (1) localized spin polarization that breaks the time-reversal symmetry and generates giant exchange spin splitting of bands; (2) a gaped inverted band structure in one spin channel (-like states above -like states), and the gaped normal band arrangement in the other spin channel – to achieve this situation a proper ratio between - and - exchange integrals is necessary; (3) the Fermi level in the gap region. That pioneering theory of the QAHE considered (Hg,Mn)Te QW as a system in which that phenomenon could be observed Liu et al. (2008), albeit in the presence of an external magnetic field, as interactions between Mn spins are antiferromagnetic according to our and previous studies Mycielski et al. (1984); Śliwa et al. (2021). However, recent experimental findings have indicated that the QHE dominates in (Hg,Mn)Te and becomes visible already in 50 mT Shamim et al. (2020), the fact linked to the resonant character of acceptor states and the formation of bound magnetic polarons Dietl (2023a, b). Under these conditions, (Hg,Mn)Te looks as a promising QHE resistance standard working in surprisingly low magnetic fields, though its performance above 50 mK is to be verified experimentally. Our first-principles results for (Hg,V)Te point to the presence of a FM coupling between V spins, an encouraging finding from the QAHE perspective. At the same time, however, our data reveal that the V impurities form a resonant donor state with the conduction band. This would mean that V doping will introduce electrons that would preclude a shift of the Fermi level toward the topological gap by a gate voltage. In the case of (Hg,Cr)Te, our ab initio studies indicate that the Cr donor level is degenerate with the valence band and, therefore, Cr acts as an isoelectronic impurity, similarly to Mn. Accordingly, it should be possible to shift the Fermi level towards the topological gap either by a suitable co-doping with electrically active impurities or by gating. At the same time, our first principles data substantiate two conclusions of the tight-binding computations Śliwa and Dietl (2023): (i) The exchange coupling between Cr spins is on the borderline between the FM case of V ions and AFM interactions between Mn ions. This opens a door for manipulation of magnetism by an electric field, strain, or pressure but suggests that the application of an external magnetic field may be necessary to polarize Cr spins. (ii) A close proximity of the Cr donor level to the top of the valence band results in a large - hybridization and, thus, in a large - exchange interaction between band carriers and localized spins, which we see as large splittings of and bands in Figs. 3 and 4. It has been shown in the companion paper Śliwa and Dietl (2023), exploiting the methodology developed earlier for HgTe and (Hg,Mn)Te QWs Novik et al. (2005); Dietl (2023b), that a large ratio between the - and - exchange integrals (denoted as and in the literature) precludes the appearance of the QAHE for the magnetization orientation along the growth direction [001] in (Hg,Cr)Te/(Cd,Hg)Te QWs. However, the spin-orbit interaction reduces the - splitting for a tilted magnetization direction. A direct calculation for the [111] magnetization direction, 3 of Hg-substitutional Cr ions in HgTe QW with a thickness of 6 nm, has demonstrated the presence of the gaped inverted band structure in one spin channel and the normal gaped band arrangement for the other spin directionŚliwa and Dietl (2023). We conclude, therefore, that (Hg,Cr)Te QWs are a perspective system for the observation of the QAHE. Furthermore, the determined magnitude of the topological energy gap of the order of 3 meV Śliwa and Dietl (2023) shows the range of Fermi level energies for which the Hall resistance will assume the quantized value Rxy= Liu et al. (2008), corresponding to the Chern number C = 1. The gate voltage width of the Hall resistance plateau will depend on the density of the in-gap localized states in a given QW Dietl (2023a).
VII Conclusions
According to theoretical studies summarized in the present and the companion paper Śliwa and Dietl (2023), the predicted magnitude and often sign of magnetic coupling between cation-substitutional Cr ions in II-VI compounds depends on the approach employed. In particular, our results show that differing conclusions about the sign of the interaction (FM vs. AFM) come out from the GGA+ approach and the hybrid functional method. This fact reflects a delicate balance between FM superexchange and mostly AFM Bloembergen-Rowland and two-electron terms Śliwa and Dietl (2023). Altogether, the data point out that while coupling between Mn ions is AFM, the FM interactions take over in the V case, but Cr in II-VI constitutes a borderline situation – FM and AFM components are of a similar strength. This may mean that the observation of the QAHE may require the application of a magnetic field to polarize Cr spins. At the same time, it might be possible to manipulate with magnetism and, thus, with topological properties by electric field, strain, pressure, and chemical content.
Another relevant question concerns the position of states brought about by magnetic impurities, which determines their electrical activity and the degree of hybridization with the band states. It has been known for a long time Dietl and Ohno (2014) that, in II-VI compounds, Mn donor and acceptor states reside deep in the valence band and high in the conduction band, respectively, so that Mn acts as an isoelectronic impurity, the fact reconfirmed by our data. Because the - interaction between Te orbitals and Hg levels, the top of the valence is high in HgTe, which results in a relatively small value of the - exchange integral in HgTe compared to other II-VI compounds Dietl and Ohno (2014); Autieri et al. (2021).
Furthermore, our results confirm the experimental finding that the V and Cr donor states are the gap of CdTe. However, according to the outcome of our computations, V states lie the conduction band of HgTe, meaning that V ions act as resonant donors. If this were the case, it would be hard to shift the Fermi level to the topological gap in (Hg,V)Te QWs. Our results indicated that the Cr donor states reside in the HgTe valence band, so that Cr ions constitute isoelectronic impurities. Since, however, those states are relatively close to the top of the valence band, we predict a relatively sizable magnitude of in (Hg,Cr)Te. According to previous Liu et al. (2008) and recent works Śliwa and Dietl (2023), a large ratio of to the - exchange integral in HgTe/(Cd,Hg)Te QWs doped with transition metals either in the QW or barriers, results in such an ordering of spin subbands, which impedes the appearance of the QAHE. Fortunately, if magnetization is tilted out of the growth direction, the spin-orbit interaction reduces the - spitting of QW states without affecting the - exchange. The presence of the QAHE is predicted for the [111] direction of magnetization in the (Hg,Cr)Te QWs Śliwa and Dietl (2023). In any case, a close energetic proximity of states and the Fermi level opens doors for a new physics in both (Hg,V)Te and (Hg,Cr)Te, not encountered in the case of (Hg,Mn)Te. A pool of other possibilities are offered by engineering superlattices or heterostructures involving TM-doped Hg-based topological systems, some of possibilities being indicated elsewhere Islam et al. (2022, 2023).
Acknowledgments
This research was supported by the Foundation for Polish Science through the IRA Programme co-financed by the EU within SG OP and EFSE (project “MagTop” no. FENG.02.01-IP.05-0028/23), by funds from the state budget allocated by the Minister of Science (Poland) as part of the Polish Metrology II programme project no. PM-II/SP/0012/2024/02. We acknowledge the access to the computing facilities of the Interdisciplinary Center of Modeling at the University of Warsaw, Grant G84-0, GB84-1, and GB84-7. We acknowledge the CINECA award under the ISCRA initiative IsC85 ”TOPMOST” and IsC93 ”RATIO” grant, for the availability of high-performance computing resources and support. We acknowledge the access to the computing facilities of the Poznan Supercomputing and Networking Center Grant No. 609.
Appendix A Transition metal doped compounds with 25% and 50% of the exact exchange
In this Appendix, we present band structures and the magnetic properties for TM-doped CdTe and HgTe obtained employing the hybrid functional with and of the exact exchange. The band structure results displayed in Figs. 6, 7, and 8 further substantiate the use of in the main body of the text. In particular, in the case of Hg0.875Mn0.125Te, the experimental value of the band gap eV Furdyna (1988) is close to the computational data for (Fig. 3) but is too small and too large for and 0.5, respectively, as shown in Figs. 7 and 8.
As we can see in Fig. 7, for we find a metallic phase in the case of HgTe doped with Cr because acceptor -states of the dopants undergo a shift below the top of the valence band. The band structures obtained within the GGA+ approach resemble the results obtained at for all compounds and, therefore, we do not display them.
In Tables 5 and Table 6 we report the values of the spin-spin couplings for the Cr-doped compounds obtained by using and 0.5, respectively. In the latter case, we find that HgTe doped with Cr shows ferromagnetic couplings.
| compound | |||
| Cd1-xVxTe | -38.1 | 22.8 | 0.6 |
| Cd1-xCrxTe | -10.5* | -16.8* | -4.8 |
| Cd1-xMnxTe | -21.9 | -21.8 | -0.5 |
| Hg1-xVxTe | 47.6 | 51.5 | -1.6 |
| Hg1-xCrxTe | -77.9 | -35.9 | -5.1 |
| Hg1-xMnxTe | -22.3 | -22.2 | -2.0 |
| compound | ||
| Cd1-xCrxTe NO JT | -218.7 | -160.3 |
| Cd1-xCrxTe JT/2 | -62.3 | 24.1 |
| Hg1-xCrxTe NO JT | 74.0 | 7.8 |
| Hg1-xCrxTe JT/2 | 88.6 | -22.7 |
Appendix B Magnetic configurations
In this Section, for completeness, we report the graphical representation of the magnetic configurations investigated in the paper. We have used 222 supercells with two magnetic impurities and we have investigated six different configurations, three in the ferromagnetic case and three in the antiferromagnetic case. The FM configurations are reported in Fig. 9, while the AFM configurations are shown in Fig. 10. In the unit cell, we have 8 cations and 8 anions. In the first configuration, the two magnetic dopants replace two cations first-neighbor, in the second configuration the magnetic dopants are second-neighbor, while in the third one, the magnetic dopants are third-neighbors. Since there are also the atoms in the repeated unit cell to be considered, (i=1,2,3) are not exactly proportional to the first-neighbor, second-neighbor and third-neighbor magnetic couplings. However, they give a strong indication on these magnetic couplings.
Appendix C Results within GGA + U approximation
In Tab. 7 we report the values obtained by using GGA + calculations. As we can see from Tab. 7, in the GGA approximation, in the case of doping with Mn, we have antiferromagnetic couplings for both HgTe and CdTe. In the case of doping with Cr and V, the couplings are ferromagnetic. The values of the third neighbor energy differences for Hg1-xCrxTe and Cd1-xCrxTe in GGA + U are unphysical. The antiferromagnetism of the couplings in the case of Mn-doped systems is still present when we use the HSE functional, therefore this property is robust, and also because already within GGA + we get the experimental insulating configuration for these systems. In GGA+ the ferromagnetism is overestimated due to the overestimation of the metallicity. In Tab. 8 we report the results obtained with GGA + U for the case of doping with Cr with JT distortions included. As in the main body of the paper, with JT/2 we indicate that we are considering the average of the positions of the atoms without and with the Jahn-Teller distortion.
| E1 | E2 | E3 | |
| Cd1-xCrxTe | 63.5 | 215.9 | 107.6 |
| Cd1-xVxTe | 497.6 | 28.2 | -1.7 |
| Cd1-xMnxTe | -15.7 | -15.6 | -0.7 |
| Hg1-xCrxTe | 168.3 | 178.4 | -36.1 |
| Hg1-xVxTe | 44.1 | 44.2 | -1.8 |
| Hg1-xMnxTe | -13.0 | -12.9 | 0.4 |
| E1 | E2 | E3 | |
| Cd1-xCrxTe NO JT | 63.5 | 215.9 | 107.6 |
| Hg1-xCrxTe NO JT | 168.3 | 178.4 | -36.1 |
| Cd1-xCrxTe JT/2 | 70.3 | 220.9 | 100.9 |
| Hg1-xCrxTe JT/2 | 171.5 | 174.4 | -32.9 |
| Cd1-xCrxTe JT | 74.8 | 140.1 | 91.4 |
| Hg1-xCrxTe JT | 171.6 | 170.5 | -30.4 |